OpenStudy (anonymous):

how do you factor poly nomials (mostly 2-3 terms) with a coefficent of x squared

OpenStudy (dumbcow):

Given ax^2 + bx + c Goal is to find factors of ac which add up to b, call them p,q p*q = ac p+q = b Example: 3x^2 +10x -8 a=3, b=10, c= -8 ac = 3*-8 = -24 what factors of -24 add up to 10 try 6 and -4 6-4 = 2 (Doesnt work) try 12 and -2 12-2 = 10 (Works) Replace 10x with 12x-2x (3x^2 +12x)+( -2x -8) 3x(x+4) -2(x+4) (x+4)(3x-2)

OpenStudy (anonymous):

Very simple one... x^2 + 2x + 1 You want the answer in the form of (x + a)(x + b). Now to find a and b. Start with the constant (1). What 2 numbers times to give +1? +1, +1, or -1, -1 Now look at the number infront of the x, what pluses to give +2 out of the above numbers? +1 and +1 therefore a = +1 and b = +1 It factorises to (x+1)(x+1). Gets slightly more complicated when there is a number infront of the x^2. Or you can just use \[-b \pm \sqrt{b^{2} - 4ac}\over2a\]